# An Explicit Entire Function of Order One with All Zeros on a Line and Bounded in a Half-Plane

### v1

@@ 8.1. First zeros of $Z_{c}$
8.2. Comparison with $\zeta$ zero statistics
The microscopic statistics of $Z_{c}$’s zeros differ fundamentally from those of $\zeta$. With 1000 zeros computed ($t\in[17.9,1420.1]$):
Statistic $\zeta$ (conditional on RH) $Z_{c}$ Clock
RMS of error term $\sqrt{\frac{\log\log T}{2\pi^{2}}}\to\infty$ $0.013$ (constant) 0
Spacing std (normalized) $\approx 0.42$ (GUE) $0.020$ 0
Pair correlation $1-\mathrm{sinc}^{2}\alpha$ peaks at integers $\delta$’s
Number variance $\Sigma^{2}(L)$ $\sim\tfrac{2}{\pi^{2}}\log L$ $\lesssim 0.25$ 0
Distribution of $S$ / $E$ Gaussian bounded, non-Gaussian 0
The zeros of $Z_{c}$ are super-rigid (nearly crystalline), with normalized spacing std two orders of magnitude below GUE. The construction captures the macroscopic transition (bounded/unbounded, AP hierarchy) but not the microscopic GUE statistics, which arise from the arithmetic content (Euler product, prime correlations) that $Z_{c}$ deliberately lacks.
@@ 8.3. Value distribution
Figure 1 shows the complex-plane images of $\log(Z_{c}/Z_{0})(\sigma+it)$, $\log\zeta(\sigma+it)$, and $\log L(\sigma+it,\chi_{4})$ as $t$ ranges over $[20,300]$, for $\sigma=1.25,1.0,0.75$.
The analytic formula gives
(10) $\log\frac{Z_{c}}{Z_{0}}(s)=-c\sum_{k=1}^{\infty}\alpha_{k}^{-s}\,e^{i\phi_{k}}+O(c),$
which has equal real and imaginary RMS: $\langle(\operatorname{Re}\log\frac{Z_{c}}{Z_{0}})^{2}\rangle=\langle(\operatorname{Im}\log\frac{Z_{c}}{Z_{0}})^{2}\rangle=\tfrac{c^{2}}{2}\sum\alpha_{k}^{-2\sigma}$. The value distribution is isotropic in the complex plane, as for $\log\zeta$ via Bohr–Jessen [3].
The quantitative difference is dramatic: at $\sigma=0.75$, the RMS of $\log(Z_{c}/Z_{0})$ is approximately $0.046$, while $\log\zeta$ has RMS $\approx 0.88$. The construction’s value distribution is confined to a bounded disk, whereas $|\log\zeta(\sigma+it)|$ is unbounded. This means $Z_{c}/Z_{0}$ cannot satisfy Voronin universality [15], and its moments do not exhibit the $(\log T)^{k^{2}}$ growth of $\zeta$. The value distribution is isotropic but concentrated—another manifestation of the missing arithmetic.
Figure 1. Complex-plane image of $\log f(\sigma+it)$ for $t\in[20,300]$. Left: $\log(Z_{c}/Z_{0})$; center: $\log\zeta$; right: $\log L(s,\chi_{4})$. Rows: $\sigma=1.25,1.0,0.75$. Color encodes $t$ (yellow==early, purple==late). Note the ${\sim}20{\times}$ difference in scale between the left column and the others.
Figure 2. Shape comparison: $\log(Z_{c}/Z_{0})$ rescaled to match the RMS of $\log\zeta$ (left) vs. $\log\zeta$ (right). The rescaled $Z_{c}$ fills a banded region (reflecting the coherent trigonometric structure) rather than the isotropic spirals of $\log\zeta$ (reflecting multiplicative independence of primes).
@@ 8.4. Numerical evidence for the linearization condition
Lemma 1.2 establishes that $\sum_{n=1}^{\infty}E(\gamma_{n}^{(0)})^{2}/n^{2}<\infty$. We verify this numerically with $K=80$ terms and $N=5000$ zero positions.
Growth of $|E(\gamma_{n}^{(0)})|$. The values $|E(\gamma_{n}^{(0)})|$ remain bounded: over the first 5000 zeros, $\max_{n}|E(\gamma_{n}^{(0)})|\approx 0.813$. A power-law regression gives $|E(\gamma_{n}^{(0)})|\sim n^{-0.025}$ (essentially flat), and a logarithmic fit yields $|E(\gamma_{n}^{(0)})|\sim(\log n)^{-0.16}$ (very slightly decreasing). The Abel-summation bound $|E(T)|=O(T)$ grossly overestimates the actual behavior at zero positions.
@@
NN $\frac{1}{N}\sum_{n=1}^{N}E(\gamma_{n}^{(0)})^{2}$ $\frac{1}{2\pi^{2}}\sum_{k=1}^{K}\alpha_{k}^{-1}$ ratio
@@
The ratio stabilizes near $1$, consistent with joint equidistribution of the $K$ phases $\{\omega_{k}\gamma_{n}^{(0)}+\phi_{k}\}$ modulo $2\pi$ (cross terms vanish, diagonal terms give the Mertens sum).
Convergence of $\sum E^{2}/n^{2}$. The partial sums of $\sum_{n=1}^{N}E(\gamma_{n}^{(0)})^{2}/n^{2}$ stabilize rapidly:
NN $\sum_{n=1}^{N}E^{2}/n^{2}$ $\sum_{n=1}^{N}\delta_{n}^{2}/\gamma_{n}^{2}$ last term $E^{2}/n^{2}$
100 0.0891 $3.79\times 10^{-5}$ $2.80\times 10^{-5}$
1000 0.0897 $88\times 10^{-5} 3.3$$30\times 10^{-7}$
5000 0.0898 $3.89\times 10^{-5}$ $3.00\times 10^{-11}$
The growth from $N=1000$ to $N=5000$ is only $0.06\%$. The linearization error $\sum\delta_{n}^{2}/\gamma_{n}^{2}\approx 3.9\times 10^{-5}$ is far below $c^{2}=2.5\times 10^{-3}$.
Spectral identity remainder. The remainder $R(t)$ in the spectral identity $\log|\Xi_{c}/\Xi_{0}|(\sigma\!+\!it)=-c\sum\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+R(t)$ has:
@@
The remainder is $O(c)$ as claimed in Proposition 3.1, and the contribution from linearization ($\sim 10^{-5}$) is negligible compared to the constant and lower-limit terms.
9. Open Problems
@@
GUE statistics. The zeros of $Z_{c}$ are super-rigid (nearly crystalline spacing, §8.2), far from the GUE statistics of $\zeta$. What additional structure on the zero perturbation $E$ would force GUE pair correlation onto the zeros of $\Xi_{c}$?
Pointwise bounds on $E$. Numerically, $|E(\gamma_{n}^{(0)})|\leq 1.05$ for all computed nn8.4), vastly smaller than the Abel bound $O(\gamma_{n})$. Proving $|E(\gamma_{n}^{(0)})|=O(1)$ would give a much tighter linearization estimate than Lemma 1.2.
Growth rate on $\sigma=1$. The Phragmén–Lindelöf argument gives $\sup_{t}|Z_{c}(1+it)|=+\infty$ but not the rate. For $\zeta$, $|\zeta(1+it)|=\Omega(\log\log t)$ [14, Thm. 8.5]. Does $|Z_{c}(1+it)|$ achieve the same rate?
10. Concluding Remarks
The function $\Xi_{c}$ provides the first explicit entire function of order $1$ that provably satisfies an analogue of the Riemann Hypothesis (all zeros on $\sigma=1/2$) together with the sharp bounded/unbounded transition at $\sigma=1$ characteristic of $\zeta$. The construction demonstrates that these properties are simultaneously achievable and provable, given sufficient control over the zero distribution.
The ratio $Z_{c}/Z_{0}$ is a Dirichlet series with generalized Euler product, providing a concrete object whose zeros are provably on the critical line and whose analytic properties (boundedness, value distribution, almost-periodicity) can be studied unconditionally. In the context of the Selberg class [12, 9], the construction shows that an RH-like zero distribution is compatible with the $\sigma=1$ transition—the two properties reinforce rather than obstruct each other.
The microscopic statistics of $Z_{c}$’s zeros differ from those of $\zeta$8.2): the zeros are nearly crystalline (super-rigid spacing) rather than GUE-distributed, and $Z_{c}$ does not satisfy Voronin universality. These differences reflect the specific arithmetic content—prime correlations, multiplicative independence of $p^{-it}$—that $\zeta$ possesses and $Z_{c}$ deliberately lacks. Understanding which additional structure would force GUE statistics onto the zeros of an RH-satisfying function remains an interesting open problem.

### v2

@@ 7.1. First zeros of $Z_{c}$
7.2. Zero statistics: theory and numerics
The microscopic statistics of $\Xi_{c}$’s zeros can be analyzed both theoretically and numerically. The key quantity is the ratio of perturbation to spacing:
(13) $\frac{|\delta_{n}|}{\text{spacing}}\approx c\,|E(\gamma_{n}^{(0)})|.$
Mean square of $E$ at the zeros. By Weyl equidistribution of the phases $\{\omega_{k}\gamma_{n}^{(0)}+\phi_{k}\}$ modulo $2\pi$ (for each finite truncation $K$, then taking $K\to\infty$):
(14) $\frac{1}{N}\sum_{n=1}^{N}E(\gamma_{n}^{(0)})^{2}\;\longrightarrow\;\frac{1}{2\pi^{2}}\sum_{k=1}^{\infty}p_{k}^{-1}\;=\;\frac{\log\log N+M}{2\pi^{2}}+o(1)$
(where $M$ is the Meissel–Mertens constant), so the RMS of $E(\gamma_{n}^{(0)})$ grows like $\sqrt{\log\log N/(2\pi^{2})}$. Consequently:
  1. (1)
    RMS of the counting error: $\bigl(\frac{1}{N}\sum(N_{c}(\gamma_{n})-n)^{2}\bigr)^{1/2}\sim c\sqrt{\log\log N/(2\pi^{2})}$, which matches the growth rate of $S(T)$ for $\zeta$ (cf. [17, Thm. 14.24]) but scaled by $c$.
  2. (2)
    Perturbation-to-spacing ratio: By (13), the RMS ratio is $c\sqrt{\log\log N/(2\pi^{2})}$, which exceeds $1$ only for $N>\exp\exp(2\pi^{2}/c^{2})$. For $c=0.05$ this threshold is $N>10^{10^{347}}$—far beyond any computation. Thus the zeros appear nearly crystalline (“super-rigid”) over any accessible range, but asymptotically the perturbations grow to many spacings, and the clock-like character is lost.
  3. (3)
    Normalized spacing variance: The deviation of the $n$-th normalized spacing from $1$ is $(\delta_{n+1}-\delta_{n})\cdot N_{\mathrm{main}}^{\prime}=c(E(\gamma_{n+1}^{(0)})-E(\gamma_{n}^{(0)}))$, which has variance $O(c^{2})$ for each $n$ (from the smoothness of $E$ on the scale of zero spacing). The normalized spacing std is therefore $O(c)$, much smaller than the GUE value $\approx 0.42$.
Numerical observations ($c=0.05$, $K=80$, 1000 zeros with $t\in[17.9,1420.1]$):
Statistic $\zeta$ (conditional on RH) $Z_{c}$ (observed)
RMS of $N-N_{\mathrm{main}}$ $\sqrt{\frac{\log\log T}{2\pi^{2}}}\to\infty$ $0.013$
Spacing std (normalized) $\approx 0.42$ (GUE [12]) $0.020$
Number variance $\Sigma^{2}(L)$ $\sim\tfrac{2}{\pi^{2}}\log L$ $\lesssim 0.25$
The observed values reflect the regime $c|E|\ll 1$ and are consistent with the theoretical $O(c)$ spacing deviation. The asymptotic growth $c\sqrt{\log\log N}$ is invisible at these heights. The construction captures the macroscopic transition (bounded/unbounded, AP hierarchy) but not the microscopic GUE statistics of $\zeta$, which arise from the Euler product and prime correlations that $Z_{c}$ deliberately lacks.
@@ 7.3. Value distribution
Figure 1 shows the complex-plane images of $\log(\Xi_{c}/\Xi_{0})(\sigma+it)$, $\log\zeta(\sigma+it)$, and $\log L(\sigma+it,\chi_{4})$ as $t$ ranges over $[20,300]$, for $\sigma=1.25,1.0,0.75$.
By the spectral identity and the $B^{2}$-AP structure of $D$, the mean square of $\log(\Xi_{c}/\Xi_{0})$ satisfies (to leading order in $c$): $\langle(\operatorname{Re}\log\frac{\Xi_{c}}{\Xi_{0}})^{2}\rangle\approx\langle(\operatorname{Im}\log\frac{\Xi_{c}}{\Xi_{0}})^{2}\rangle\approx\tfrac{c^{2}}{2}\sum p_{k}^{-2\sigma}$ (to leading order, from the $B^{2}$-norm of $D$). The value distribution is isotropic in the complex plane, as for $\log\zeta$ via Bohr–Jessen [3].
The quantitative difference is dramatic: at $\sigma=0.75$, the RMS of $\log(\Xi_{c}/\Xi_{0})$ is approximately $0.046$, while $\log\zeta$ has RMS $\approx 0.88$. The construction’s value distribution is confined to a bounded disk, whereas $|\log\zeta(\sigma+it)|$ is unbounded. This means $\Xi_{c}/\Xi_{0}$ cannot satisfy Voronin universality [18], and its moments do not exhibit the $(\log T)^{k^{2}}$ growth of $\zeta$ (cf. [11]). The value distribution is isotropic but concentrated—another manifestation of the missing arithmetic.
Figure 1. Complex-plane image of $\log f(\sigma+it)$ for $t\in[20,300]$. Left: $\log(\Xi_{c}/\Xi_{0})$; center: $\log\zeta$; right: $\log L(s,\chi_{4})$. Rows: $\sigma=1.25,1.0,0.75$. Color encodes $t$ (yellow==early, purple==late). Note the ${\sim}20{\times}$ difference in scale between the left column and the others.
Figure 2. Shape comparison: $\log(\Xi_{c}/\Xi_{0})$ rescaled to match the RMS of $\log\zeta$ (left) vs. $\log\zeta$ (right). The rescaled curve fills a banded region (reflecting the coherent trigonometric structure) rather than the isotropic spirals of $\log\zeta$ (reflecting multiplicative independence of primes).
@@ 7.4. Numerical evidence for the linearization condition
Lemma 3.2 establishes that $\sum_{n=1}^{\infty}E(\gamma_{n}^{(0)})^{2}/n^{2}<\infty$. We verify key aspects numerically with $K=80$ terms and $N=5000$ zero positions.
@@
NN $\frac{1}{N}\sum_{n=1}^{N}E(\gamma_{n}^{(0)})^{2}$ $\frac{1}{2\pi^{2}}\sum_{k=1}^{K}p_{k}^{-1}$ ratio
@@
The ratio stabilizes near $1$, consistent with joint equidistribution of the $K$ phases $\{\omega_{k}\gamma_{n}^{(0)}+\phi_{k}\}$ modulo $2\pi$.
Spectral identity remainder. The remainder $R(t)$ in the spectral identity $\operatorname{Re}\log(\Xi_{c}/\Xi_{0})(\sigma\!+\!it)=-c\sum p_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+R(t)$ has:
@@
The remainder is $O(c)$ as claimed, and the linearization contribution ($\sim 10^{-5}$) is negligible compared to the constant and lower-limit terms.
8. Open Problems and Concluding Remarks
The function $\Xi_{c}$ provides an explicit entire function of order $1$ that provably satisfies an analogue of RH (all zeros on $\sigma=1/2$) together with the sharp bounded/unbounded transition at $\sigma=1$ characteristic of $\zeta$. The ratio $\Xi_{c}/\Xi_{0}$ is, to leading order, a Dirichlet series with generalized Euler product (Proposition 5.1), providing a concrete object whose zeros are provably on the critical line and whose analytic properties can be studied unconditionally.
The construction demonstrates that the $\sigma=1$ transition is driven entirely by the marginal convergence/divergence of the Dirichlet series $D(s)=\sum e^{-ik\theta}p_{k}^{-s}$ at $\sigma=1$, in direct analogy with $\log\zeta(s)=P(s)+O(1)$ [17, §3.2]. In the context of the Selberg class, this shows that an RH-like zero distribution is compatible with the $\sigma=1$ transition—the two properties reinforce rather than obstruct each other.
@@ 8. Open Problems and Concluding Remarks
GUE statistics. The zeros of $Z_{c}$ are super-rigid (nearly crystalline spacing, §7.2), far from the GUE statistics of $\zeta$. What additional structure on the perturbation would force GUE pair correlation onto the zeros of $\Xi_{c}$?
Pointwise bounds on $E$. The mean square (14) shows that $|E(\gamma_{n}^{(0)})|\sim\sqrt{\log\log n}$ in RMS, but the worst-case bound (2) gives only $O(\gamma_{n}^{1/2+\varepsilon})$. Numerically, $|E(\gamma_{n}^{(0)})|\leq 1.05$ for all computed nn7.4). Proving that $|E(\gamma_{n}^{(0)})|=O(\sqrt{\log\log n})$ pointwise (matching the RMS rate) would sharpen the linearization estimate and clarify when the perturbations first exceed the zero spacing.
Growth rate on $\sigma=1$. The Phragmén–Lindelöf argument gives $\sup_{t}|Z_{c}(1+it)|=+\infty$ but not the rate. For $\zeta$, $|\zeta(1+it)|=\Omega(\log\log t)$ [17, Thm. 8.5]. Does $|Z_{c}(1+it)|$ achieve the same rate?
An actual Dirichlet series. The ratio $\Xi_{c}/\Xi_{0}$ is approximately a Dirichlet series with Euler product (Proposition 5.1), but the bounded correction factor $M(s)$ prevents it from being one exactly. The base function $\Xi_{0}$ is not a Dirichlet series either (its Fourier exponents $\log\gamma_{n}^{(0)}$ are not of the form $\log n$). Can one unconditionally construct a Dirichlet series $\sum a_{n}\,n^{-s}$ (convergent for $\sigma>1/2$) with all zeros provably on $\sigma=1/2$ and the $\sigma=1$ transition? The Davenport–Heilbronn example [4] shows that a functional equation alone does not suffice; some form of multiplicative structure appears necessary.
Acknowledgments
This paper was developed through a collaboration between the author and Claude (Anthropic). The author posed the original question (whether one can construct an explicit entire function with the Riemann Hypothesis and the $\sigma=1$ transition), proposed the unit-disc analogy that motivated the Hadamard product approach, identified that a family result is insufficient (insisting on a single function), and formulated the equidistribution criterion for unboundedness. Claude proposed the Hadamard product construction, developed the spectral identity and its proof, found the dyadic large-sieve argument that resolved the linearization convergence, suggested the $\alpha_{k}=p_{k}$ switch that yields the Dirichlet series structure, and wrote the paper.